OpenStudy (anonymous):

1. What is the first step in solving g-3/7 >5?(Points : 1) Add 3 to each side. Subtract 3 from each side. Divide each side by 7. Multiply each side by 7. 2. Write an inequality to represent the problem (1 pt.) and then solve the inequality by writing the pairs which solve it (1 pt.). Find all sets of two consecutive positive odd integers whose sum is less than or equal to 18. (Points : 2)

OpenStudy (anonymous):

is it \(\frac{g-3}{7}>5\) ?

OpenStudy (anonymous):

Get rid of fractions immediately. ie Multiply by 7. Always your best option

OpenStudy (anonymous):

hello!!

OpenStudy (anonymous):

@satellite73 It doesn't really matter anyway. It's a fraction so you should multiply by 7.

OpenStudy (anonymous):

good point

OpenStudy (anonymous):

actually try \(n+n+2\leq 18\)

OpenStudy (anonymous):

this tell you \(2n+2\leq 18\) \[2n\leq 16\] \[n\leq 8\]so you don't have too many pairs of odd integers to write

OpenStudy (anonymous):

You should probably set a lower bound: 0 \[0 < n + (n + 2) \le 18\]\[n \in \mathbb{Z}\]